OpenStudy (anonymous):

5/7 - ( 3/14 + 3/14) Use any method to add or subtract. Help Please?!

OpenStudy (anonymous):

8/7

OpenStudy (anonymous):

its 8/7

OpenStudy (dayakar):

simplify the bracket terms first 3/14 +3/14 = 6/14 =2/7

OpenStudy (anonymous):

the answers are A) 2/7 B) 1/21 C) 1/7 D)1 1/7

OpenStudy (anonymous):

Thank You! @dayakar

OpenStudy (dayakar):

5/7-2/7 = 3/7

OpenStudy (dayakar):

welcom :)

OpenStudy (anonymous):

are you able to help he more? @dayakar

OpenStudy (anonymous):

me**

OpenStudy (dayakar):

i will try

OpenStudy (anonymous):

Jerome has 1/4 of the groups video games at his house. Mario has 2/5 of the groups video games at his house. What fraction of the groups video games is either at Jerome's house of Mario's house?

OpenStudy (anonymous):

@dayakar

OpenStudy (dayakar):

is there any options

OpenStudy (dayakar):

do u have any idea

OpenStudy (anonymous):

do you add it? @dayakar

OpenStudy (dayakar):

2/5*1/4 = 1/5

OpenStudy (anonymous):

Rebecca had 9/10 of a yard of ribbon. Then she gave 7/20 of a yard to a friend. How much does she have left? This is the last one @dayakar